Equation: N2+3H2->2NH3
What is the chemical excess in moles?
1
Expert's answer
2017-05-16T06:42:36-0400
Solution:
1 mol N2 reacts with 3 moles H2
If 1 mol H2 reacts with 1 mol N2
1 mol N2 ‒ 3 moles H2
x mol N2 ‒ 1 mol H2
x = 1/3 mol = 0.3 mol N2
N2 is in excess, n(excess) = 1 – 0.3 = 0.7 mol N2